Solve Problems Using Linear Systems
15. A motor boat took 5h to travel a distance of 60 km up a river, against the current. The return trip took 3h. Find the average speed of the boat in still water and the speed of the current.
Thank you sooooooo much!!!!

Answer by scott8148(6628)

"A motor boat took 5h to travel a distance of 60 km up a river" ___ so b-c=60/5 or 12

"The return trip took 3h" ___ so b+c=60/3 or 20

adding equations ___ (b-c)+(b+c) = 12+20 ___ 2b = 32 ___ b=16

substituting __ (16)+c = 20 ___ c = 4